# Subsystem Thermalization Hypothesis in Quantum Spin Chains with Conserved Charges

### v1

@@ IV Thermalization demographics of quantum spin-$1/2$ XXZ chains with $U(1)$ symmetry
To consider spin chains with more rich symmetries, such as non-Abelian ones, we consider the following $L$-site spin-$1/2$ XXZ spin chain model with the following Hamiltonian [36]
$H=J\sum_{i=1}^{L}(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}+\Delta S_{i}^{z}S_{i+1}^{z})$
$+\lambda J\sum_{i=1}^{L}(S_{i}^{x}S_{i+2}^{x}+S_{i}^{y}S_{i+2}^{y}+\Delta S_{i}^{z}S_{i+2}^{z})$
@@
[…]
@@
    $H$ commutes with $\vec{S}_{\rm tot}:=\sum_{k=1}^{L}\vec{S}_{k}$ if $d=h=0$ and $\Delta=1$. This special model with non-Abelian $SU(2)$ symmetry is called the XXX model.
@@
    This model is chaotic if the NNN coupling $10<\lambda<$ 1 and $h=0$ (no random disorder) for both XXZ ($\Delta\neq 0$) and XX ($\Delta=0$) models. This can be verified by fitting the level spacing statics to the Wigner-Dyson distribution, as shown in Fig. 19 of Appendix A.
@@
Figure 12: Thermalization demographics of the population ratios of $S^{a}_{A}$ for the XXZ chain of (22) with $U(1)$ conserved charge for p-GGE specified only by $\beta_{1}$ for various almost-superselection states with with fixed $\langle Q_{0}\rangle=$ (a) $0.0$, (b) $0.2$, (c) $0.4$, (d) $0.6$, and (e) $0.8$, respectively. This is the fine-grained version of Fig. 21 of Appendix A. The thermalization hypothesis works well for $A=1$ (blue) but not so for $A=2$ (red) and $A=3$ (green).
@@
We consider the two conserved charges of this model denoted by
@@
In the next section, we will consider the XXX model with $\vec{S}_{\rm tot}$ as the $SU(2)$ conserved charges. Since we have studied the subsystem thermalization hypothesis for the Ising chain with $\Pi$ as the $Z_{2}$ conserved charge, we will ignore this discrete charge for the following considerations.
As in the cases of quantum Ising chains, we consider the demographics of relative entropies $S_{A}^{t,a,e}$ for both GGE and p-GGEs. The demographics of $S_{A}^{t}$ bear similar patterns as in the Ising cases; that is, the subsystem thermalization hypothesis holds well for $A=1$ and worsens as $A$ increases. Thus, we will not show the results in the main text but in Appendix A.
As for the thermalization demographics of $S_{A}^{a}$, we find something interesting when compared with the Ising counterparts. In the Ising case, the subsystem thermalization hypothesis works well for $A=1$ for all fixed values of $\langle Q_{1}\rangle$, but only for small values of $\langle Q_{0}\rangle$. However, as shown in Fig. 11 and 12, we see the role of $Q_{0}$ and $Q_{1}$ swap, that is, the subsystem thermalization hypothesis works well for $A=1$ for all fixed values of $\langle Q_{0}\rangle$, but only for small values of $\langle Q_{1}\rangle$.
@@
Finally, we consider the subsystem ETH for the XXZ chains. With loss of generality, we only compare with the GGE states specified by $(\beta_{0},\beta_{1})$. The result is shown in Fig. 13. We see the subsystem ETH holds well for $A=1$ but worsens as $A$ increases. Also, the plot patterns for the XXZ chain of $U(1)$ charges differ from the Ising ones with $Z_{2}$ charges.
@@ V.1 Demographics of typical-state thermalization
We first consider the subsystem thermalization hypothesis for the typical states compared to GGE states (or NATS), specified by the chemical potentials $\beta_{\mu=0,1,2,3}$. It is impossible to simultaneously express the demographics of $S^{t}_{A}$ of four chemical potentials. Instead, we look into the dependence of each chemical potential of the full demographics. The results are shown in Fig. 14, in which we also show the demographics of $\langle Q_{\mu=0,1,2,3}\rangle$ of the typical states. Again, the subsystem thermalization holds well for $A=1$ and worsens as $A$ increases. Our result confirms that NATS/GGE states can be adopted for the thermalization hypothesis of the typical states even though the conserved charges are noncommuting  [^2] Our result, in fact, confirms the non-Abelian ETH proposed in [26] based on the subsystem version of the thermalization hypothesis for typical states. This is because the ETH considered in [26] is examined by comparing the time average of the expectation values of the local observables on the evolved pure states with the corresponding thermal average over the NATS/GGE ensemble states. The former ones are equivalent to the average of the expectation values of the local observables on the typical states..
@@
We now move to the p-GGEs. To demonstrate the richness of the p-GGEs, we first consider the subsystem thermalization hypothesis by comparing the typical states with the p-GGE states with (i) single conserved charges, (ii) two conserved charges, and (iii) three conserved charges. The results are presented in Fig. 15, 16 and 17, respectively. The overall features of these results are as follows. First, the subsystem thermalization hypothesis works better for the p-GGEs with more conserved charges. Second, for the p-GGEs of the same number of charges, the subsystem thermalization hypothesis works better for the ones involving no $Q_{0}$. Thus, if we restrict to $A=1$, our results suggest that the typical-state thermalization hypothesis holds quite generically for the p-GGEs with at least two conserved charges without involving $Q_{0}$. This implies that NATSs are not the only thermal ensemble states adopted for the subsystem thermalization hypothesis for the typical states. Indeed, our framework with p-GGEs extends the thermalization hypothesis to a more general scope. In particular, it is quite interesting to see that the thermalization hypothesis also works for those p-GGEs with $Q_{0}=H$ excluded. This implies that the Hamiltonian may not be essential for considering quantum thermalization. This is also implied by the fact that the (inverse) temperatures for some of the corresponding GGE or p-GGE are negative, which bears no physical meaning in realistic thermodynamics. This could be related to the thermalization hypothesis working better for the p-GGEs without $Q_{0}$ than those with $Q_{0}$.
@@ V.2 Demographics of eigenstate thermalization
Finally, we consider the subsystem ETH for the XXX chains with $SU(2)$ conserved charges. Since the $SU(2)$ charges are noncommuting, we choose the eigenstates to be specified by the energy $E$ and the $z$-componet of the total spin, i.e., $S^{z}_{\rm tot}$. For such eigenstates, the expectation values of $S^{x}_{\rm tot}$ and $S^{y}_{\rm tot}$ vanishes, as expected by the usual argument for the Stern-Gerlach experiment. Thus, it is only sensible to consider subsystem ETH with the p-GGEs specified by either $\beta_{0}$ or $\beta_{3}$ or the GGE specified by both. As usual, the subsystem thermalization hypothesis holds well for $A=1$ and worsens as $A$ increases. Thus, we only show the demographics of $S_{A=1}^{e}$ in Fig. 18. The results show that the subsystem ETH works almost equally well for GGE and p-GGE specified by $\beta_{3}$, both better than the p-GGE specified by $\beta_{0}$. This is similar to the previous cases for typical states and for the XXZ chain.
@@ VI Conclusion
Quantum thermalization is fascinating because it yields deep implications for the second-law perspective of pure states. Compared to the thermalization for the open quantum system or classical system, quantum thermalization can manifest even without taking the thermodynamic limit. This implies that we can implement the exact diagonalization method to examine the thermalization hypothesis of a not-so-large system as long as the subsystem size is small enough compared to the system size. This provides the starting point for the numerical plots done in this work. The usual considerations for the thermalization hypothesis are done for the systems without conserved charges so that the corresponding thermal ensemble for comparison is either the micro-canonical or canonical ensemble. This work extensively considers the subsystem thermalization hypothesis of the typical states or eigenstates for the systems with conserved charges of $Z_{2}$, $U(1)$ or $SU(2)$ symmetries. Furthermore, we generalize the thermal ensembles beyond the generalized Gibbs ensemble (GGE). We call these thermal ensembles the partial-GGE (p-GGE), of which some conserved charges are excluded. Moreover, in the framework of p-GGEs, the Hamiltonian and other conserved charges are treated on equal footing. Moreover, our results show that the p-GGEs with $H$ excluded usually yield better results for quantum thermalization than the ones with $H$.
Based on the framework of the subsystem thermalization hypothesis with p-GGEs, we can quantify the validity of this scheme by the smallness of the relative entropy between the subsystem’s reduced states obtained from the typical states/eigenstates and the corresponding p-GGEs. We can then examine the universality of quantum thermalization with various p-GGEs by numerically calculating the demographics of relative entropies. Our results of thermalization demographics demonstrate that the subsystem thermalization hypothesis of the typical states (and their fine-grained versions) and eigenstates compared to p-GGEs holds well quite generically as long as the subsystem size is small enough. However, there are some cases where the subsystem typical-state thermalization hypothesis fails, for example, for the p-GGEs of $SU(2)$ case with one or two conserved charges, especially with the Hamiltonian included. Despite that, our results show that the thermalization hypothesis can be extended to p-GGEs even without including Hamiltonian. This greatly enlarges the scope of quantum thermalization and will deserve future studies for more general systems with conserved charges and larger sizes.
@@ Acknowledgements.
Appendix A Typical-state thermalization demographics of the XXZ chain with U(1) conserved charges
@@
As mentioned in the main text, This model is chaotic if the NNN coupling $10<\lambda<$ 1 and $h=0$ (no random disorder) for both XXZ ($\Delta\neq 0$) and XX ($\Delta=0$) models. We verify this by fitting the level spacing statics to the Wigner-Dyson distribution, as shown in Fig. 19. This is consistent with the thermalization hypothesis. For readers’ interest, we also present the density plots of three quantities to characterize the GGE thermal states associated with the same numerically model considered in Fig. 19.
@@
We now consider the demographics of the thermalization hypothesis. As we have presented the plots for $S^{a,e}_{A}$ in the main text, we will only show the ones for the typical states. Although we consider the $U(1)$ charge, not the $Z_{2}$ of this XXZ model for the GGE or p-GGE states, the typical-state thermalization demographics bear similar patterns and features as in the Ising case with only $Z_{Z}$ conserved charge. The XXZ model we consider is described by the Hamiltonian (IV) with the coupling parameters given by $L=10$, $\Delta=\lambda=0.5$, $h=d=0$, and $J=1$. The demographics for GGE specified by $(\beta_{0},\beta_{1})$, and p-GGE specified by $\beta_{0}$ and $\beta_{1}$ are given in the Fig. 20, 21 and 22, respectively. The thermalization hypothesis seems to work better in Fig. 22 than in Fig. 21.
@@
Figure 20: Density plots of typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $(\beta_{0}$, $\beta_{1})$ for the XXZ chain of (22) when comparing with the GGE states specified by $(\beta_{0},\beta_{1})$, i.e., respectively the inverse temperature and the chemical potential to the $Q_{1}=S^{z}_{\rm tot}$. The thermalization hypothesis works well for $A=1$ (a) and worsens as $A$ increases, i.e., for $A=3$ (b) and $A=5$ (c).
@@
Figure 21: (a) Typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $\beta_{0}$ for the XXZ chain of (22) by comparing with a p-GGE specified only by $\beta_{0}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, i.e., $A=3$ (blue circles) and $A=5$ (black triangles). (b) Demographics of the typical states classified by the $\langle Q_{0}\rangle$.
@@
Figure 22: (a) Typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $\beta_{1}$ for the XXZ spin chain of (22) by comparing with a p-GGE specified only by $\beta_{1}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, i.e., $A=3$ (blue circles) and $A=5$ (black triangles). (b) Demographics of the typical states classified by the $\langle Q_{0}\rangle$.
@@
The result shows that the subsystem typical-state thermalization hypothesis works well for $A=1$ and worsens as $A$ increases, as in the Ising cases. Despite that, it is interesting to compare the demographic patterns of Fig. 20 with its Ising counterpart shown in Fig. 4. We see that the current patterns are more uniform on the $(\beta_{0},\beta_{1})$-plane due to the continuous symmetry instead of the discrete one.
@@
The fine-grained versions of Fig. 21 and Fig. 22 as the corresponding almost-superselection-state thermalization demographics shown in Fig. 12 and Fig. 11, respectively. We see that the almost-superselection-state thermalization hypothesis for $A=1$ works well for the p-GGE specified by $\beta_{1}$, not for the p-GGE specified by $\beta_{0}$ for which the thermalization hypothesis worsens as $\langle Q_{1}\rangle$ increases away from zero. Interestingly, the almost-superselection-state thermalization demographics of quantum Ising chains, as shown in Fig. 7 and 8, works for the p-GGE specified by $\beta_{0}$, but not for the p-GGE specified by $\beta_{1}$. It is the opposite for the ZZX chains.

### v2

@@ IV Thermalization demographics of quantum spin-$1/2$ XXZ chains with $U(1)$ symmetry
To consider spin chains with richer symmetries, such as non-Abelian ones, we consider the following $L$-site spin-$1/2$ XXZ spin chain model with the following Hamiltonian [65]
$H=J\sum_{i=1}^{L-1}(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}+\Delta S_{i}^{z}S_{i+1}^{z})$
$+\lambda J\sum_{i=1}^{L-2}(S_{i}^{x}S_{i+2}^{x}+S_{i}^{y}S_{i+2}^{y}+\Delta S_{i}^{z}S_{i+2}^{z})$
@@
where $\vec{S}_{i}:=\vec{\sigma}_{i}/2$ is the spin vector operator at the $i$-th site. Besides, the parameter $dJ$ denotes the Zeeman splitting at the defect site. The Zeeman splitting $h_{i}J$ reflects onsite disorder due to random static magnetic fields, with $h_{i}$ being random values drawn from a uniform distribution within [h,h][-h,h], where $h$ indicates the disorder strength. The anisotropy parameter $\Delta$ determines the interaction type: the system is isotropic when $\Delta=1$, meaning the Ising interaction and the flip-flop term have equal strength. The parameter $\lambda$ characterizes the relative size of the nearest-neighbor (NN) coupling to the next-nearest-neighbor (NNN) coupling.
@@
    $H$ commutes with $\vec{S}_{\rm tot}:=\sum_{k=1}^{L}\vec{S}_{k}$ if $d=h=0$ and $\Delta=1$. This special model, characterized by a non-Abelian $SU(2)$ symmetry, is known as the XXX model.
@@
    This model is chaotic if the NNN coupling $0<\lambda<1$ and $h=0$ (no random disorder) for both XXZ ($\Delta\neq 0$) and XX ($\Delta=0$) models. This can be verified by fitting the level spacing statistics to the Wigner-Dyson distribution, as shown in Fig. 22 of Appendix A.
@@
Figure 12: Thermalization demographics of the population ratios of $S^{a}_{A}$ for the XXZ chain of (22) with $U(1)$ conserved charge for p-GGE specified only by $\beta_{1}$ for various almost-superselection states with fixed $\langle Q_{0}\rangle=$ (a) $0.0$, (b) $0.2$, (c) $0.4$, (d) $0.6$, and (e) $0.8$, respectively. This is the fine-grained version of Fig. 24 of Appendix A. The thermalization hypothesis works well for $A=1$ (blue) but not so for $A=2$ (red) and $A=3$ (green).
@@
We consider the two conserved charges of this model, denoted by
@@
For completeness, in Fig. 26 of Appendix C, we also present the results with p-GGE, including $Q_{2}=\Pi$, which demonstrates that the thermalization hypothesis remains effective, as in the Ising chains. In the next section, we will consider the XXX model with $\vec{S}_{\rm tot}$ as the $SU(2)$ conserved charges. Since we have studied the subsystem thermalization hypothesis for the Ising chain with $\Pi$ as the $Z_{2}$ conserved charge, we will ignore this discrete charge for the following considerations.
As in the cases of quantum Ising chains, we consider the demographics of relative entropies $S_{A}^{t,a,e}$ for both GGE and p-GGEs. The demographics of $S_{A}^{t}$ exhibit similar patterns to those in the Ising cases; that is, the subsystem thermalization hypothesis holds well for $A=1$ and worsens as $A$ increases. Thus, we will not present the results in the main text, but rather in Appendix A.
Regarding the thermalization demographics of $S_{A}^{a}$, we observe an interesting difference when compared with its Ising counterparts. In the Ising case, the subsystem thermalization hypothesis works well for $A=1$ for all fixed values of $\langle Q_{1}\rangle$, but only for small values of $\langle Q_{0}\rangle$. However, as shown in Fig. 11 and 12, we see the role of $Q_{0}$ and $Q_{1}$ swap, that is, the subsystem thermalization hypothesis works well for $A=1$ for all fixed values of $\langle Q_{0}\rangle$, but only for small values of $\langle Q_{1}\rangle$.
@@
Finally, we consider the subsystem ETH for the XXZ chains. With some loss of generality, we only compare with the GGE states specified by $(\beta_{0},\beta_{1})$. The result is shown in Fig. 13. We observe that the subsystem ETH performs well for $A=1$ but deteriorates as $A$ increases. Additionally, the plot patterns for the XXZ chain of $U(1)$ charges differ from those for the Ising chain with $Z_{2}$ charges.
@@ V.1 Demographics of typical-state thermalization
We first consider the subsystem thermalization hypothesis for the typical states compared to GGE states (or NATS), specified by the chemical potentials $\beta_{\mu=0,1,2,3}$. It is impossible to simultaneously express the demographics of $S^{t}_{A}$ of four chemical potentials. Instead, we look into the dependence of each chemical potential on the full demographics. The results are shown in Fig. 14, in which we also show the demographics of $\langle Q_{\mu=0,1,2,3}\rangle$ of the typical states. Again, the subsystem thermalization holds well for $A=1$ and worsens as $A$ increases. Our result confirms that NATS/GGE states can be adopted for the thermalization hypothesis of the typical states, even though the conserved charges are noncommuting  [^3] Our result, in fact, confirms the non-Abelian ETH proposed in [47] based on the subsystem version of the thermalization hypothesis for typical states. This is because the ETH considered in [47] is examined by comparing the time average of the expectation values of the local observables on the evolved pure states with the corresponding thermal average over the NATS/GGE ensemble states. The former ones are equivalent to the average of the expectation values of the local observables on the typical states..
@@
We now move to the p-GGEs. To demonstrate the richness of the p-GGEs, we first consider the subsystem thermalization hypothesis by comparing the typical states with the p-GGE states with (i) single conserved charges, (ii) two conserved charges, and (iii) three conserved charges. The results are presented in Fig. 15, 16 and 17, respectively. The overall features of these results are as follows. First, the subsystem thermalization hypothesis is more effective for p-GGEs with more conserved charges. Second, for the p-GGEs of the same number of charges, the subsystem thermalization hypothesis works better for the ones involving no $Q_{0}$. Thus, if we restrict to $A=1$, our results suggest that the typical-state thermalization hypothesis holds quite generically for the p-GGEs with at least two conserved charges without involving $Q_{0}$. This implies that NATSs are not the only thermal ensemble states adopted for the subsystem thermalization hypothesis for the typical states. Indeed, our framework with p-GGEs extends the thermalization hypothesis to a more general scope. In particular, it is quite interesting to observe that the thermalization hypothesis also applies to those p-GGEs with $Q_{0}=H$ excluded. This implies that the Hamiltonian may not be essential for considering quantum thermalization. This is also implied by the fact that the (inverse) temperatures for some of the corresponding GGE or p-GGE are negative, which bears no physical meaning in realistic thermodynamics. This could be related to the thermalization hypothesis working better for the p-GGEs without $Q_{0}$ than those with $Q_{0}$.
@@ V.2 Demographics of eigenstate thermalization
Finally, we consider the subsystem ETH for the XXX chains with $SU(2)$ conserved charges. Since the $SU(2)$ charges are noncommuting, we choose the eigenstates to be specified by the energy $E$ and the $z$-component of the total spin, i.e., $S^{z}_{\rm tot}$. For such eigenstates, the expectation values of $S^{x}_{\rm tot}$ and $S^{y}_{\rm tot}$ vanish, as expected by the usual argument for the Stern-Gerlach experiment. Thus, it is only sensible to consider subsystem ETH with the p-GGEs specified by either $\beta_{0}$ or $\beta_{3}$ or the GGE specified by both. As usual, the subsystem thermalization hypothesis holds well for $A=1$ and worsens as $A$ increases. Thus, we only show the demographics of $S_{A=1}^{e}$ in Fig. 18. The results show that the subsystem ETH performs almost equally well for both GGE and p-GGE specified by $\beta_{3}$, which is better than the p-GGE specified by $\beta_{0}$. This is similar to the previous cases for typical states and the XXZ chain.
@@ VI Conclusion
Quantum thermalization is fascinating because it yields deep implications for the second-law perspective of pure states. Compared to thermalization in open quantum systems or classical systems, quantum thermalization can manifest even without taking the thermodynamic limit. This implies that we can implement the exact diagonalization method to examine the thermalization hypothesis of a not-so-large system, as long as the subsystem size is sufficiently small compared to the system size. This provides the starting point for the numerical plots done in this work. The usual considerations for the thermalization hypothesis are applied to systems without conserved charges, so that the corresponding thermal ensemble for comparison is either the microcanonical or canonical ensemble. This work extensively considers the subsystem thermalization hypothesis for typical states or eigenstates of systems with conserved charges of $Z_{2}$, $U(1)$, or $SU(2)$ symmetries. Furthermore, we generalize the thermal ensembles beyond the generalized Gibbs ensemble (GGE). We refer to these thermal ensembles as the partial-GGE (p-GGE), from which some conserved charges are excluded. Moreover, in the framework of p-GGEs, the Hamiltonian and other conserved charges are treated on equal footing. Thus, when considering the thermalization hypothesis, one can decide which information to discard by choosing the corresponding p-GGE.
Based on the framework of the subsystem thermalization hypothesis with p-GGEs, we can quantify the validity of this scheme by the smallness of the relative entropy between the subsystem’s reduced states obtained from the typical states/eigenstates and the corresponding p-GGEs. We can then examine the universality of quantum thermalization with various p-GGEs by numerically calculating the demographics of relative entropies. Our results on thermalization demographics demonstrate that the subsystem thermalization hypothesis, which applies to typical states (and their fine-grained versions) and eigenstates compared to p-GGEs, holds quite generally as long as the subsystem size is small enough. We have also performed a size-scaling study in Appendix B to demonstrate that the thermalization hypothesis is more effective for larger systems.
However, there are some cases where the subsystem typical-state thermalization hypothesis fails, for example, in the p-GGEs of the $SU(2)$ case with one or two conserved charges, especially when the Hamiltonian is included. Moreover, in some cases. e.g., comparing Fig. 11 and Fig. 12, or see Fig. 15 and Fig.16, we observe that the p-GGEs with $H$ excluded yield sharper demographic around $S^{t}_{A}\simeq 0$ than the ones with $H$. In these cases, the thermalization hypothesis works better for the $H$-excluded p-GGEs than the ones with $H$ included. This indicates that the thermalization hypothesis can be extended to p-GGEs even without including the Hamiltonian. This significantly expands the scope of quantum thermalization and warrants future studies for more general systems with conserved charges and larger sizes.
@@ Acknowledgements.
Appendix A Size-scaling study of typical-state thermalization of the quantum Ising chain
Figure 19: Size-scaling of typical-state thermalization demographics of the population ratios of $S^{t}_{A=1,2,3}$ for different lattice sizes $L=6,8,10$ of a GGE state for the Ising chain of (19) with conserved charges $Q_{0,1}$ given by (20).
Figure 20: Size-scaling of typical-state thermalization demographics of the population ratios of $S^{t}_{A=1,2,3}$ for different lattice sizes $L=6,8,10$ of a p-GGE state for the Ising chain of (19) by including only the conserved charges $Q_{0}=H$.
Figure 21: Size-scaling of typical-state thermalization demographics of the population ratios of $S^{t}_{A=1,2,3}$ for different lattice sizes $L=6,8,10$ of a p-GGE state for the Ising chain of (19) by including only conserved charges $Q_{1}=\Pi$.
Fig. 19 shows the relative entropy results for different subsystem sizes $A=1,2,3$ and lattice sizes $L=6,8,10$ in the 1D Quantum Ising model. The simulations are performed with transverse and longitudinal fields $h_{x}=1.05$, $h_{z}=0.5$, and consider two conserved symmetries: the Hamiltonian $Q_{0}=H$ and the parity operator $Q_{1}=\Pi$. Similar to the fig. 19, Fig. 20 considers $Q_{0}=H$ , while Fig. 21 considers $Q_{1}=\Pi$ only. In Fig. 19  20  21(a), it shows the relative entropy probability distributions for the size of the subsystem $A=1$, comparing the three sizes of the system. Fig. 19  20  21(b) and (c) extend this comparison to $A=2$ and $A=3$, respectively.
As the total system size $L$ increases, the distributions of relative entropy become sharper and more peaked, indicating a more well-defined thermal behavior. For smaller systems $(L=6)$, the distributions of the relative entropy are significantly broader. In contrast, for $L=10$, the relative entropy distributions display a more concentrated profile, consistent with stronger thermalization. This observation indicates that larger systems exhibit more pronounced thermalization behavior in their subsystems. Thermalization becomes more evident as the lattice size increases, which is consistent with the eigenstate thermalization hypothesis framework applied to integrable and near-integrable systems.
Appendix B Typical-state thermalization of the XXZ chain with $U(1)$ conserved charges
@@
As mentioned in the main text, this model is chaotic if the NNN coupling $0<\lambda<1$ and $h=0$ (no random disorder) for both XXZ ($\Delta\neq 0$) and XX ($\Delta=0$) models. We verify this by fitting the level spacing statistics to the Wigner-Dyson distribution, as shown in Fig. 22. This is consistent with the thermalization hypothesis. For readers’ interest, we also present the density plots of three quantities to characterize the GGE thermal states associated with the same numerically model considered in Fig. 22.
@@
We now consider the demographics of the thermalization hypothesis. As we have presented the plots for $S^{a,e}_{A}$ in the main text, we will only show the ones for the typical states. Although we consider the $U(1)$ charge, rather than the $Z_{2}$ of this XXZ model, for the GGE or p-GGE states, the typical-state thermalization demographics exhibit similar patterns and features to those in the Ising case, with only $Z_{Z}$ conserved charge. The XXZ model we consider is described by the Hamiltonian (IV) with the coupling parameters given by $L=10$, $\Delta=\lambda=0.5$, $h=d=0$, and $J=1$. The demographics for GGE specified by $(\beta_{0},\beta_{1})$, and p-GGE specified by $\beta_{0}$ and $\beta_{1}$ are given in the Fig. 23, 24 and 25, respectively. The thermalization hypothesis seems to work better in Fig. 25 than in Fig. 24.
@@
Figure 23: Density plots of typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ vs. $(\beta_{0}$, $\beta_{1})$ for the XXZ chain of (22) when comparing with the GGE states specified by $(\beta_{0},\beta_{1})$, i.e., respectively the inverse temperature and the chemical potential to the $Q_{1}=S^{z}_{\rm tot}$. The thermalization hypothesis works well for $A=1$ (a) but worsens as $A$ increases, as seen for $A=3$ (b) and $A=5$ (c).
@@
Figure 24: (a) Typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $\beta_{0}$ for the XXZ chain of (22) by comparing with a p-GGE specified only by $\beta_{0}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, as seen for $A=3$ (blue circles) and $A=5$ (black triangles). (b) Demographics of the typical states classified by the $\langle Q_{0}\rangle$.
@@
Figure 25: (a) Typical-state thermalization demographics: $S^{t}_{A=1,3,5}$ v.s. $\beta_{1}$ for the XXZ spin chain of (22) by comparing with a p-GGE specified only by $\beta_{1}$. The thermalization hypothesis also works well for $A=1$ (red squares), but it worsens as $A$ increases, as seen for $A=3$ (blue circles) and $A=5$ (black triangles). (b) Demographics of the typical states classified by the $\langle Q_{0}\rangle$.
@@
The results show that the subsystem typical-state thermalization hypothesis works well for $A=1$ and worsens as $A$ increases, similar to the Ising cases. Despite that, it is interesting to compare the demographic patterns of Fig. 23 with its Ising counterpart shown in Fig. 4. We see that the current patterns are more uniform on the $(\beta_{0},\beta_{1})$-plane due to the continuous symmetry instead of the discrete one.
@@
The fine-grained versions of Fig. 24 and Fig. 25 as the corresponding almost-superselection-state thermalization demographics shown in Fig. 12 and Fig. 11, respectively. We see that the almost-superselection-state thermalization hypothesis for $A=1$ works well for the p-GGE specified by $\beta_{1}$, not for the p-GGE specified by $\beta_{0}$, for which the thermalization hypothesis worsens as $\langle Q_{1}\rangle$ increases away from zero. Interestingly, the almost-superselection-state thermalization demographics of quantum Ising chains, as shown in Fig. 7 and 8, work for the p-GGE specified by $\beta_{0}$, but not for the p-GGE specified by $\beta_{1}$. It is the opposite for the ZZX chains.
Appendix C Typical-state thermalization of the XXZ chain with parity symmetry
Figure 26: Typical-state thermalization demographics of the population ratios of $S^{t}_{A=1,2,3}$ for the XXZ chain of (22) for the p-GGE of three, two, and one charges (Here $Q_{0}=H$, $Q_{1}=S^{z}_{\rm tot}$, $Q_{2}=\Pi$.): (a) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{0},\beta_{1},\beta_{2})$, (b) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{0})$, (c) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{1})$, (d) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{2})$, (e) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{0},\beta_{1})$, (f) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{1},\beta_{2})$, and (g) $S^{t}_{A=1,2,3}$ v.s. $(\beta_{0},\beta_{2})$.
Fig. 26 shows the relative entropy results for different subsystem sizes $A=1,2,3$ in the 1D Quantum XXZ model with lattice length $L=10$. The system is analyzed under the parameter conditions $\Delta=\lambda=0.5$ and $J=1$, taking into account three conserved charges denoted by $Q_{0}=H,\;\;Q_{1}=S_{total}^{z},$ and $Q_{2}=\Pi$. We consider the subsystem thermalization hypothesis for typical states and energy eigenstates by comparing them locally with GGE states and various types of p-GGE states.
Relative entropy is used here as a diagnostic tool to examine the thermalization behavior of subsystems. It measures the distinguishability between the actual reduced density matrix of a subsystem and its corresponding thermal (or generalized Gibbs ensemble) reference state. A lower relative entropy suggests that the subsystem is effectively thermalized, whereas higher values indicate deviations from thermal equilibrium.